Find the coordinates of the x and y intercepts. x + 2y = 4
Answer by MathTutorer(13)

Using the formula y=mx+c
x + 2y = 4 -x on both sides we get 2y = -x + 4
divide both sides by 2
then you get
y intercept is when x = 0
x intercept is when y = 0
y = -1/2x + 2
y = 2
Now time for x
x + 2y = 4
subtract 2y to both sides
x = -2y + 4
x = 4
Gradient = -1/2
Hope that helps ya :D